Institut for Matematik og Datalogi Odense Universitet

{May 16, 1995 \atop JFB}

DM15 – F95 – Week 30

Lecture, May 15

We covered up through section 2.5.5 in the second chapter of Mitchell’s monograph.

Lecture, May 22

We will do a review of the course.

Discussion section, May 29

This will be the 15th discussion section this semester. Peter Høyer will lead it, in U6, starting at 10:15 AM.

Question and answer session, June 7

Peter Høyer will lead a question and answer session, as review for the exam, starting at 10:15 AM, in U47.